OpenStudy (anonymous):

if you make random quesses for a 10 multiple choice test with 5 possible answers each what is the probabilit of getting at least 1 correct?

OpenStudy (raden):

P(at least 1 correct) = 1 - P(all question are false)

OpenStudy (raden):

and you have : P(correct) = 1/5 and P(false) = 4/5

OpenStudy (raden):

so, it is P(at least 1 correct) = 1 - (4/5)^10

OpenStudy (anonymous):

thank you